# Post-Quantum Cryptography (PQC) Network Instrument: Measuring PQC Adoption Rates and Identifying Migration Pathways

### v1

@@ II Approach Overview
This section describes the network topology, architecture of the network optical tap, placement of our network instrument, data collected, and statistical methods used.
Figure 1 describes our network instrument. Using FABRIC testbed (A), we deployed our instrument at the NCSA site (B) to ingest Zeek connection metadata and parse session and application layer cipher suite information (C, D, E). An example of PQC key exchange and statistics that we parsed is shown for Secure Shell (SSH) protocol (F, G). Our instrument is embedded in a nation-scale network for monitoring a wide spectrum of scientific workloads as follows.
@@
Ethical considerations. Our approach works on connection metadata produced by Zeek parsers. We did not use man-in-the-middle or traditional SSL termination equipment. Thus, we only see the encrypted data and not the original content. However, we do have access to the timing information of connection establishment to derive the trend and seasonality of the PQC adoption rate. The connection metadata generally includes standard information such as IP address, certificate, and negotiation of cipher suites. We do not see Personally Identifiable Information (PII).
@@ VIII Lessons Learned &
End-to-end PQC requires both client-server support. To completely support PQC, we have found that both client and server have to support the same cipher suite, and widespread adoption hinges upon effective protocol negotiation when establishing the initial handshake, followed by key exchanges and digital signatures. For effective post-quantum security to be possible, it is thus necessary that as many machines as possible adopt post-quantum versions of network protocols.
@@
Remark 9: In the future, we will release an open dataset containing metadata of cryptographic suites across major network protocols (SSH, TLS, RDP, etc.) upon the paper’s publication. With our network instrument in place, we will track the adoption of PQC algorithms, give feedback to NIST, expect to discover potential novel PQC attacks in the wild and help migrate HPC applications to be quantum-safe. Finally, we will expand our network instrument to a set of networked vantage points (PQC telescopes) to continuously measure PQC adoption at the Internet scale.
@@ ACKNOWLEDGEMENTS
This work was partly supported by the National Science Foundation under contract CCF #2319190. We acknowledge Dr. Jim Basney for the useful discussions about SciTokens, Dr. Santiago Nunez-Corrles, Dr. Edoardo Giusto, members of NCSA Quantum Task Force, Dependable Classical-Quantum Computing Systems Engineering Working Group, and the Illinois Quantum Information Science and Technology Center (IQUIST) faculty members for insightful feedback. We also want to recognize the following organizations/programs: NCSA Student Pushing Innovation Program, IBM-Illinois Discovery Accelerator Institute, FABRIC, PPoSS, and the National Science Foundation’s CyberCorps Scholarship for Service (SFS) Program, NCSA Integrated Cyberinfrastructure/IRST team, and NIST HPC Security Working Group. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of their employers or the sponsors.

### v2

@@ II Approach Overview
This section describes the network topology, architecture of the network optical tap, placement of our network instrument, data collected, and statistical methods used. Figure 1 describes our network instrument. Using FABRIC testbed (A), we deployed our instrument at the NCSA site (B) to ingest Zeek connection metadata and parse session and application layer cipher suite information (C, D, E). An example of PQC key exchange and statistics that we parsed is shown for the Secure Shell (SSH) protocol (F, G). Our instrument is embedded in a nation-scale network for monitoring a wide spectrum of scientific workloads as follows.
@@
Both NCSA and FABRIC provide a wide gamut of HPC workloads that we will analyze for quantum-resistant cryptography.
@@
Ethical considerations. Our approach works on connection metadata produced by Zeek. We did not use man-in-the-middle or traditional SSL termination equipment. Thus, we only see the encrypted data and not the original content. However, we do have access to the timing information of connection establishment to derive the trend and seasonality of the PQC adoption rate. The connection metadata generally includes standard information such as IP address, certificate, and negotiation of cipher suites. We do not see Personally Identifiable Information (PII).
@@ VIII Lessons Learned &
End-to-end PQC requires both client-server support. To completely support PQC, we have found that both client and server have to support the same cipher suite. Widespread adoption hinges upon effective protocol negotiation when establishing the initial handshake, followed by key exchanges and digital signatures. For effective post-quantum security to be possible, it is thus necessary that as many machines as possible adopt post-quantum versions of network protocols.
@@
Remark 9: We are releasing an open dataset containing metadata of cryptographic suites across major network protocols (SSH, TLS, RDP, etc.) upon the paper’s publication at https://pmcao.github.io/pqc. With our network instrument in place, we will track the adoption of PQC algorithms, give feedback to NIST, expect to discover potential novel PQC attacks in the wild, and help migrate HPC applications to be quantum-safe. Finally, we will expand our network instrument to a set of networked vantage points (PQC telescopes) to continuously measure PQC adoption at the Internet scale.
@@ ACKNOWLEDGEMENTS
We acknowledge Dr. Jim Basney for in-depth discussions about SciTokens, Dr. Santiago Nunez-Corrales, Dr. Edoardo Giusto, members of NCSA Quantum Task Force, Dependable Classical-Quantum Computing Systems Engineering Working Group, and the Illinois Quantum Information Science and Technology Center (IQUIST) faculty members, particularly Dr. Brian DeMarco, for insightful feedback. This work was partly supported by the National Science Foundation (NSF) under contract CCF #2319190. We also want to recognize the following organizations/programs: the NSF’s Trusted CI Cybersecurity Center of Excellence, NCSA Student Pushing Innovation Program (SPIN), Illinois Computes, IBM-Illinois Discovery Accelerator Institute, FABRIC, PPoSS, and the NSF’s CyberCorps Scholarship for Service (SFS) Program, ACCESS and Delta allocations, NCSA Integrated Cyberinfrastructure/IRST team, particularly Timothy Boerner, James Eyrich, Ryan Walker, Christopher Clausen, and Dr. Yang Guo at the NIST HPC Security Working Group. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of their employers or the sponsors.